### v3

@@ Results
CT-CHAT generates clinically accurate and detailed answers to questions about the given CT image, as shown in Figure 6b. To evaluate the performance, we employ multiple metrics, including BLEU-1, METEOR, ROUGE-L, and CIDEr, which assess the quality of the generated text. Additionally, inspired by previous work [27], we utilize an LLM-based metric, named the Llama score, to evaluate the clinical accuracy of responses, with the models’ outputs compared against ground truths (see benchmarking the CT-CHAT model section of Methods). The results indicate that CT-CHAT, being capable of evaluating the full information of 3D CT volumes, consistently outperforms the other open-source vision-language AI assistants across all evaluated tasks. It shows superior ability in generating long and short answers, handling multiple-choice questions, and producing detailed radiology reports that align closely with clinical expectations, as shown in Figure 6c. Extended Data Figure 6 provides a detailed breakdown, demonstrating that CT-CHAT significantly outperforms others in all tasks.
There are two key reasons why CT-CHAT significantly outperforms 2D-based models: (1) Image: CT-CHAT leverages the 3D spatial information processed by CT-CLIP’s pretrained vision transformer, which captures much richer and more detailed anatomical structures than 2D models—a critical advantage for CT interpretation. (2) Text: CT-CHAT is trained on 3D CT volumes paired with radiology-specific linguistic data from CT-RATE, enabling it to generate highly precise and clinically relevant responses, far surpassing those produced by models trained on general-purpose texts. This powerful combination of CT-CLIP’s capacity to handle complex 3D imaging data and the domain-specific paired image-text data from CT-RATE allows CT-CHAT to push the boundaries of multimodal tasks in 3D medical imaging. In summary, CT-CHAT not only surpasses existing 2D tools but also provides a scalable solution for enhancing chest CT interpretation, underscoring the need for generalist models in 3D medical imaging.
In the absence of existing vision-language baselines for 3D medical imaging, we extend our evaluation of CT-CHAT beyond comparisons with 2D multimodal AI assistants by benchmarking its report generation capabilities against state-of-the-art 3D radiology report generation models—specifically RadFM [29] and CT2Rep [30]. This task is particularly important as it closely reflects the daily workflow of radiologists and provides a more clinically relevant benchmark than VQA. Additionally, the language generation metrics used in our VQA evaluation (such as BLEU and ROUGE) are primarily text-focused and may introduce bias due to differences in model training objectives. They are also limited in assessing the clinical fidelity of responses. Therefore, to robustly evaluate report generation quality, we employ clinical efficacy metrics (precision, recall, and F1 score), which better reflect the diagnostic value of generated reports. We extract these labels using our automated report labeler—the same model used to annotate CT-RATE—and benchmark CT-CHAT using its 8B Meta Llama 3.1 configuration. Notably, even though the RAD-ChestCT dataset lacks text reports, it includes binary abnormality labels, allowing us to evaluate generated reports by applying our text labeler and computing metrics accordingly.
CT-CHAT consistently outperforms both RadFM and CT2Rep on the CT-RATE internal and RAD-ChestCT external validation sets across all three clinical metrics, as shown in Figure 6e. The only exception is precision on the internal set, where RadFM slightly outperforms CT-CHAT. However, this is attributed to RadFM generating fewer positive findings, leading to low false positives—but also to correspondingly low recall and F1 scores—indicating limited clinical utility.
The superior performance of CT-CHAT over CT2Rep is especially noteworthy, given that both models are developed using the CT-RATE dataset. Unlike CT2Rep, which does not leverage pretraining, CT-CHAT benefits from CT-CLIP’s pretrained vision encoder, highlighting the strength of our foundational model. These results reinforce the effectiveness of CT-CLIP in enabling high-quality downstream applications such as clinically meaningful report generation and demonstrate the value of leveraging contrastively learned representations in 3D medical imaging.
@@ Discussion
Radiology reports provide comprehensive descriptions of medical images without adding extra labeling burden, as they are already part of the daily routine for radiologists [48]. Recent studies show that foundational models, trained on paired radiology reports and 2D medical images, perform comparably to fully supervised methods in various tasks [15]. However, these studies have not extended to 3D medical images, primarily because a publicly available 3D medical imaging dataset paired with text reports is absent. Therefore, CT-RATE, which pairs 3D chest CT volumes with radiology reports (Figure 1a), represents a significant advancement as the first of its kind. In this study, we introduce CT-RATE, a novel and comprehensive dataset, which we first utilize to develop CT-CLIP (Figure 1c). To our knowledge, CT-CLIP is the first foundation model in 3D medical imaging utilizing radiology text reports.
@@
CT-CLIP provides a versatile solution for various tasks without requiring task-specific training. Unlike supervised methods, it eliminates the need for explicit labels and enables predictions beyond predefined classes (Figure 1e), making it particularly valuable when learning objectives evolve post-training. CT-CLIP’s zero-shot multi-abnormality detection outperforms the state-of-the-art supervised method on the internal validation set across all metrics (Figure 4a). Importantly, this experiment suggests that explicit labels are not necessary for high performance in 3D medical image interpretation when corresponding radiology reports are available for training. Furthermore, we externally validate CT-CLIP on two independent datasets, Rad-ChestCT and UPMC, from a different country (Figure 3e). The model’s ability to generalize to different data distributions, a primary challenge in deploying medical AI [49], is confirmed as it outperforms the fully supervised method in all metrics in the external validation sets as well (Figure 4a). To establish a fully supervised baseline and fine-tune CT-CLIP, we extract multi-abnormality labels from radiology reports, making them publicly available. Following recent examples, we develop an automatic label extractor by finetuning a language model (Figure 1b) [50]. However, developing such a system requires significant time, expertise, and domain knowledge, as evidenced by our effort in manually annotating a small portion of the reports. CT-CLIP, as a zero-shot classifier, overcomes this bottleneck by detecting abnormalities at a supervised level without the need for an automatic labeler. Finetuning experiments further confirm CT-CLIP’s representational learning capabilities, demonstrating superior performance (Figure 4a). Despite being trained without explicit supervision, CT-CLIP’s vision encoder captures clinically meaningful features that are highly transferable to downstream tasks. These finetuning strategies—one tailored for classification with predefined labels (ClassFine), and the other enabling open-vocabulary inference (VocabFine)—build upon the same pretrained encoder and consistently yield superior results compared to a fully supervised baseline. This indicates that the encoder does not rely on memorizing specific abnormality classes but instead learns a more general and semantically aligned representation of the 3D chest CT volumes. Such flexibility underscores the foundational nature of CT-CLIP and its potential for broad applicability in medical imaging tasks, even beyond those explicitly considered during training. Finally, and indicating a direction for future developments, the effective clustering observed in t-SNE projections (Figure 4d-e) unveils another aspect of CT-CLIP’s clinical potential. The successful clustering of abnormalities (Extended Data Figures 1 and 2) suggests that CT-CLIP could extend its utility beyond multi-abnormality detection to assist in prognosis, disease grading, and the calculation of lung-related parameters in future research.
@@
Analyzing CT-CLIP’s open-vocabulary inference capabilities, we observe significant potential for developing multimodal algorithms that address broader and more diverse clinical applications. To expand these capabilities, we finetune four state-of-the-art LLMs through visual instruction tuning with the CT tokens created by CT-CLIP, resulting in the creation of CT-CHAT. For finetuning, we create and open-source a visual question-answering dataset derived from CT-RATE (Extended Data Figure 8). As a result, CT-CHAT significantly outperforms existing 2D models, underscoring the urgent need for generalist models specialized in 3D medical imaging (Figure 6c). We find the choice of LLM (Supplementary Table 10a) and the use of concept alignment pretraining (Supplementary Table 10b) have minimal impact on CT-CHAT’s accuracy. In contrast, including the volume as input is crucial, particularly in report generation (Supplementary Table 10b). This highlights the value of the CT-RATE dataset and CT-CLIP in developing a generalist AI assistant for 3D medical imaging. Also, we demonstrate that engineering the system prompt of CT-CHAT (Supplementary Table 9) helps mitigate pathology hallucinations when the volume is not provided (Supplementary Figure 11). Furthermore, our model, with new special tokens, is capable of responding in various modes, including long and short answers, multiple-choice, and report generation (Extended Data Figure 9). To further evaluate CT-CHAT’s clinical utility, we benchmark its report generation performance against state-of-the-art 3D medical models RadFM [29] and CT2Rep [30]. Since radiology report generation closely mirrors the daily workflow of radiologists, it serves as a more clinically relevant benchmark than VQA. Instead of relying solely on conventional text similarity metrics—which are limited in assessing the clinical accuracy of generated content—we employ clinical efficacy metrics, using our in-house automated report labeler, which was also used to annotate CT-RATE (see creating the CT-RATE dataset section of Methods for more details). CT-CHAT consistently outperforms both RadFM and CT2Rep on the CT-RATE internal and RAD-ChestCT external validation sets across all clinical metrics (Figure 6e). The only exception is precision on the internal set, where RadFM scores slightly higher due to generating much fewer positive findings. Notably, CT-CHAT’s advantage over CT2Rep—despite both models being trained on CT-RATE—highlights the effectiveness of pretraining with CT-CLIP. These results reinforce the strength of our foundational model, CT-CLIP, in learning clinically meaningful representations for 3D medical imaging and enabling high-quality performance in downstream tasks. That said, despite CT-CHAT setting a new state-of-the-art in report generation for 3D computed tomography, it remains far from being ready for clinical deployment. As with all AI systems, challenges such as robustness, interpretability, regulatory compliance, and prospective validation must be addressed before integration into routine radiology workflows.To help broaden accessibility, we introduce a user-friendly graphical interface (also open-sourced via our GitHub repository) that allows non-experts to interact with CT-CHAT effortlessly. This interface not only makes CT-CHAT a valuable tool for educational and clinical settings but also invites a wider audience to engage with and contribute to the evolution of interactive AI-driven inference in clinical practice.
@@
Despite these advancements, our CT-RATE dataset and the CT-CLIP and CT-CHAT models have certain limitations. First, biases inherent to our single-institution dataset may influence the performance of our self-supervised training. Although we curate a large and diverse set of volume-report pairs, the data originates from a single geographic and clinical setting, which may limit the generalizability of our models to other populations or institutions. We conduct an ablation study to assess the impact of training set size on CT-CLIP’s performance, revealing that performance increases with larger training sets (Figure 4c). This suggests that incorporating data from multiple institutions and more diverse patient populations could further enhance model robustness and generalizability. Second, CT-CLIP depends on the specific terminology used in reports and the choice of prompts during zero-shot detection. Variability in how diseases are described across radiology reports can affect detection accuracy. To assess the importance of prompt engineering, we evaluate zero-shot multi-abnormality detection performance using seven different prompts during inference (Figure 4b). While we select the best-performing prompt from commonly used structures in radiology reports, continued optimization of prompts may further improve performance. Third, the evaluation of CT-CHAT faces limitations due to the lack of pretrained 3D medical multimodal AI assistants for benchmarking. As a result, we compare CT-CHAT against 2D models, which may not fully capture the 3D spatial information inherent in CT volumes. To address this, we utilize digitally reconstructed radiographs (DRRs) generated from our dataset [47]. This approach is suggested by prior work showing that models trained with DRRs can yield comparable results to those trained with 3D chest CT volumes [47]. However, the absence of true 3D benchmarks highlights the need for further development and validation of multimodal AI assistants in 3D medical imaging. Due to the absence of 3D VQA baselines, we also benchmark CT-CHAT’s report generation function against state-of-the-art 3D radiology report generation models. Despite CT-CHAT being trained for VQA, CT-CHAT outperforms these models; however, the quality of its generated reports remains insufficient for clinical use, underscoring the need for further improvements in performance and clinical validation. Finally, our current models are limited to chest CT volumes. To broaden their clinical applicability, future work should focus on extending CT-RATE, CT-CLIP, and CT-CHAT to include other 3D imaging modalities such as MRI and PET scans, as well as other anatomical regions. Expanding to different modalities and body parts would enhance the utility of our models across a wider range of medical imaging tasks and clinical scenarios.
In conclusion, the development of CT-CLIP and CT-CHAT marks a significant advancement in 3D medical imaging, particularly for interpreting chest CT volumes. Enabled by the creation of CT-RATE, the first open-source 3D imaging dataset paired with radiology reports, these models set a benchmark for multimodal AI in 3D medical imaging. The open-source availability of CT-RATE, CT-CLIP, and CT-CHAT is expected to be a valuable asset to the medical AI community, providing a solid foundation for future research. These foundational models also can be deployed in hospitals with minimal finetuning, making them especially useful in emergency settings where radiologists may not always be available, thereby enhancing diagnostic capabilities and knowledge dissemination in 3D medical imaging.
@@
Figure 1: Figure 1: Overview of our dataset and CT-CLIP. a. Data acquisition protocol for the CT-RATE dataset, comprising non-contrast 3D chest CT volumes and corresponding radiology text reports written by expert radiologists. b. Extraction of multi-abnormality labels from radiology reports using a finetuned text encoder. c. Training of our CT-CLIP framework, which pairs chest CT volumes with radiology reports from the CT-RATE dataset through a contrastive learning approach. d. Illustration of the contrastive learning procedure, employed in the correlation matrix during training to ensure precise match recognition and mismatch differentiation. $\text{V}_{\text{i}}$ shows volume embeddings and $\text{R}_{\text{i}}$ shows report embeddings. e. Inference for zero-shot multi-abnormality detection utilizes normalized probabilities derived from softmax-applied contrastive loss for each positive and negative prompt.
@@
Figure 2: Figure 2: Comprehensive analysis of the novel CT-RATE dataset. a. CT-RATE comprises data from 21,304 unique patients, with ages ranging from 18 to 102 years. b. For validation and finetuning purposes, multi-abnormality labels for 18 distinct abnormalities are extracted from the corresponding radiology reports for each CT volume. c. The training process of the CT-CLIP model utilizes the impression and findings sections of radiology reports, which vary in length, along with the chest CT volumes. d. The number of slices in the chest CT volumes within CT-RATE ranges from 100 to 600. e. Spacings in the x-axis, y-axis, and z-axis vary across volumes; all chest CT volumes are preprocessed to ensure consistent spacing before training each model. f. Despite originating from a single institution, CT-RATE includes volumes acquired using scanners from three different manufacturers, ensuring variability.
@@
Figure 4: Figure 4: A detailed evaluation of CT-CLIP a. Multi-abnormality classification performance is evaluated on an internal and two external validation sets using four key metrics. Our CT-CLIP outperforms the fully supervised baseline across all metrics. Performance improves with both finetuning techniques, demonstrating the robustness of the CT-CLIP’s encoder in feature extraction from CT volumes. The scores are color-highlighted to emphasize the performance increase compared to random benchmarks. b.Different prompts are utilized for the zero-shot multi-abnormality detection task with CT-CLIP. The best-performing prompt, highlighting the importance of prompt engineering, is selected for further experiments. c. To explore how the performance scales with dataset size, we train CT-CLIP using different portions of CT-RATE. (9.8% of CT-RATE matches the size of the only currently available open-source chest CT dataset with multi-abnormality labels.) Contour lines denote $\pm$1 standard deviations calculated via bootstrapping. d. t-SNE projections map the distribution of the number of abnormalities present in cases, utilizing the latent values from individual reports and CT volumes in CT-RATE. e. t-SNE projections illustrate the age distribution of patients in the dataset. The clustering shows that CT volumes and reports from older patients tend to correspond with a higher number of reported abnormalities, indicating a correlation between age and disease prevalence.
@@
Figure 5: Figure 5: Utilizing CT-CLIP for retrieval of chest CT volumes a. CT-CLIP can be used for volume-to-volume retrieval utilizing the vision encoder within CT-CLIP. b. CT-CLIP can be used for report-to-volume retrieval utilizing both the vision and text transformers within CT-CLIP. c. Comprehensive evaluation of volume-to-volume retrieval performance on internal and external validation sets, with scores color-highlighted to emphasize improvements over random scores, shows similar retrieval accuracy even for the finetuned models. d. The evaluation of report-to-volume retrieval is limited to the internal validation set due to the absence of paired radiology reports in the external validation set. Notably, finetuning CT-CLIP for detection diminishes its retrieval performance. This reduction in performance is primarily observed in report-to-volume retrieval, indicating that the VocabFine finetuning predominantly impacts on the text transformer component of CT-CLIP.
@@
Figure 6: Figure 6: Overview and detailed evaluation of CT-CHAT a. The architecture of CT-CHAT, a vision-language AI assistant for 3D chest CT volumes, trained by finetuning LoRA weights on a state-of-the-art large language model using CT tokens encoded by the 3D vision encoder of CT-CLIP. b. An illustration showing an example CT-CHAT output, highlighting its superior performance as a vision-language assistant specifically designed for 3D chest CT volumes compared to its 2D counterparts with DRRs provided. c. The question-answering performance of various vision-language assistants is evaluated across five key metrics, with radar charts providing a detailed breakdown of evaluations for each of the four different visual question-answering tasks. d. An ablation study on the accuracy of CT-CHAT using different LLMs, shows that CT-CHAT with the Llama 3.1 model performs slightly better. e. Report generation performance is evaluated using clinical efficacy metrics, calculated by adapting CT-RATE’s automatic report labeler to assess clinical correctness across multiple findings on two test sets.

### v4

@@ Results
CT-CHAT generates clinically accurate answers to questions about the given CT, as in Figure 6b. To evaluate, we employ multiple metrics, including BLEU, METEOR, ROUGE-L, and CIDEr, which assess the quality of the generated text. Also, inspired by previous work [27], we use an LLM-based metric, Llama score, to evaluate the clinical accuracy of responses, with the models’ outputs compared against ground truths (see benchmarking the CT-CHAT model section of Methods). The results indicate that CT-CHAT, being capable of evaluating the full information of 3D volumes, consistently outperforms the other open-source vision-language AI assistants across all tasks. It shows superior ability in generating long and short answers, handling multiple-choice questions, and producing detailed radiology reports that align closely with clinical expectations, as shown in Figure 6c. Extended Data Figure 6 provides a detailed breakdown, demonstrating that CT-CHAT significantly outperforms others in all tasks. There are two key reasons why CT-CHAT significantly outperforms 2D-based models: (1) Image: CT-CHAT leverages the 3D spatial information processed by CT-CLIP’s pretrained vision transformer, which captures much richer and more detailed anatomical structures than 2D models, a critical advantage for CT interpretation. (2) Text: CT-CHAT is trained on 3D CT volumes paired with radiology-specific linguistic data from CT-RATE, enabling it to generate highly precise and clinically relevant responses, far surpassing those produced by models trained on general-purpose texts. This powerful combination of CT-CLIP’s capacity to handle complex 3D imaging data and the domain-specific paired image-text data from CT-RATE allows CT-CHAT to push the boundaries of multimodal tasks in 3D medical imaging. In summary, CT-CHAT not only surpasses existing 2D tools but also provides a scalable solution for enhancing chest CT interpretation, underscoring the need for generalist models in 3D medical imaging.
In the absence of existing vision-language baselines for 3D medical imaging, we extend our evaluation of CT-CHAT beyond comparisons with 2D multimodal AI assistants by benchmarking its report generation capabilities against state-of-the-art 3D radiology report generation models, specifically RadFM [29] and CT2Rep [30]. This task is particularly important as it closely reflects the daily workflow of radiologists and provides a more clinically relevant benchmark than VQA. To robustly assess report generation accuracy, we conducted expert evaluations with three board-certified radiologists who independently rated 150 reports for each model (CT-CHAT, RadFM, and CT2Rep). To ensure a fair and comprehensive assessment, we did not select cases randomly; instead, we used the Llama score on CT-CHAT outputs to stratify cases into low (0–3), medium (4–6), and high (7–10) clinical accuracy ranges, selecting 50 cases from each range. Each radiologist then scored the reports on a scale from 0 to 10 based on predefined clinical criteria. The mean expert scores demonstrate CT-CHAT’s clear superiority: Radiologist 1 rated CT-CHAT at 3.44 (vs. RadFM 0.55, CT2Rep 2.66); Radiologist 2 at 4.10 (vs. RadFM 0.91, CT2Rep 2.52); and Radiologist 3 at 4.52 (vs. RadFM 0.86, CT2Rep 2.74). These results are presented in Figure 6d.
To further strengthen report generation, we enhanced CT-CHAT with structured 3D segmentation inputs by providing detailed nodule information,such as lobar location, centroid, diameters, and volumes, extracted using a state-of-the-art 3D nodule segmentation algorithm on the CT-RATE training set. For implementation details on how we integrate nodule segmentation information, see the developing the CT-CHAT model section of Methods. This additional structured input significantly improved report accuracy, as shown in Figure 6e, demonstrating that leveraging high-quality 3D segmentation models meaningfully boosts clinical performance. Currently, we incorporate only nodule information; incorporating structured details for other abnormalities would likely enhance performance further, but robust 3D segmentation models for these additional findings are not yet available.
For quantitative evaluation, we exclude language generation metrics because they are limited in capturing clinical fidelity. Instead, we use clinical efficacy metrics (precision, recall, F1, and CRG score) that better reflect diagnostic value. Labels are extracted with our automated report labeler, the same model used to annotate CT-RATE, and CT-CHAT is benchmarked using its 8B Meta Llama 3.1 configuration. Although the RAD-ChestCT dataset lacks text reports, its binary abnormality labels allow us to apply the text labeler and compute metrics consistently. CT-CHAT outperforms RadFM and CT2Rep on the CT-RATE internal and RAD-ChestCT external validation sets across all metrics, as shown in Figure 6e. CT-CHAT’s superior performance over CT2Rep is particularly notable, given that both models are trained on CT-RATE. Unlike CT2Rep, which does not leverage pretraining, CT-CHAT benefits from CT-CLIP’s pretrained vision encoder, demonstrating the strength of our foundational approach. Our results show that integrating outputs from other advanced 3D algorithms, such as robust segmentation tools, can further improve tasks like report generation. These findings highlight that continued development and integration of high-quality 3D models are essential for clinically deployable AI systems. By providing first-of-their-kind resources and benchmarks, especially through CT-RATE, our work takes an important step toward this goal, though significant challenges remain for routine clinical use.
@@ Discussion
Radiology reports provide comprehensive descriptions of medical images without adding extra labeling burden, as they are already part of the daily routine for radiologists [47]. Recent studies show that foundational models, trained on paired radiology reports and 2D medical images, perform comparably to fully supervised methods in various tasks [15]. However, these studies have not extended to 3D medical images, primarily because a publicly available 3D medical imaging dataset paired with text reports is absent. In this study, we introduce CT-RATE, a comprehensive dataset, which we first use to develop CT-CLIP, a foundation model for 3D medical imaging using radiology text reports (Figure 1c).
@@
CT-CLIP provides a versatile solution for various tasks without requiring task-specific training. Unlike supervised methods, it eliminates the need for explicit labels and enables predictions beyond predefined classes (Figure 1e), making it particularly valuable when learning objectives evolve post-training. CT-CLIP’s zero-shot multi-abnormality detection outperforms the state-of-the-art supervised method across all metrics (Figure 4a). Importantly, this experiment suggests that explicit labels are not necessary for high performance in 3D medical image interpretation when corresponding radiology reports are available for training. Furthermore, we externally validate CT-CLIP on two independent datasets, Rad-ChestCT and UPMC, from a different country (Figure 3e). The model’s ability to generalize to different distributions, a primary challenge in deploying medical AI [48], is confirmed as it outperforms the fully supervised method in all metrics in the external validation sets as well (Figure 4a). To establish a fully supervised baseline and finetune CT-CLIP, we extract multi-abnormality labels from reports, making them publicly available. Following recent examples, we develop an automatic labeler by finetuning a language model (Figure 1b) [49]. However, developing such a system requires time, expertise, and domain knowledge, as evidenced by our effort in manually annotating a small portion of the reports. CT-CLIP, as a zero-shot classifier, overcomes this bottleneck by detecting abnormalities at a supervised level without the need for an automatic labeler. Finetuning experiments further confirm CT-CLIP’s representational learning capabilities, demonstrating superior performance (Figure 4a). Despite being trained without explicit supervision, CT-CLIP’s vision encoder captures clinically meaningful features that are highly transferable to downstream tasks. These finetuning strategies, one tailored for classification with predefined labels (ClassFine), and the other enabling open-vocabulary inference (VocabFine), build upon the same pretrained encoder and consistently yield superior results compared to a fully supervised baseline. This indicates that the encoder does not rely on memorizing specific abnormality classes but instead learns a more general and semantically aligned representation of the 3D chest CT volumes. Such flexibility underscores the foundational nature of CT-CLIP and its potential for broad applicability in medical imaging tasks, even beyond those explicitly considered during training. Finally, and indicating a direction for future developments, the effective clustering observed in t-SNE projections (Figure 4d-e) unveils another aspect of CT-CLIP’s clinical potential. The successful clustering of abnormalities (Extended Data Figures 1 and 2) suggests that CT-CLIP could extend its use beyond multi-abnormality detection to assist in prognosis, disease grading, and the calculation of lung-related parameters in future research.
@@
Analyzing CT-CLIP’s open-vocabulary inference capabilities, we observe potential for developing multimodal algorithms that address broader and more diverse clinical applications. To expand these capabilities, we finetune four state-of-the-art LLMs through visual instruction tuning with the CT tokens created by CT-CLIP, resulting in the creation of CT-CHAT. For finetuning, we create and open-source a visual question-answering dataset derived from CT-RATE (Extended Data Figure 8). As a result, CT-CHAT outperforms existing 2D models, underscoring the urgent need for generalist models specialized in 3D medical imaging (Figure 6c). We find that the choice of LLM (Supplementary Table 10a) and the use of concept alignment pretraining (Supplementary Table 10b) have minimal impact on CT-CHAT’s accuracy. In contrast, including the volume as input is crucial, particularly in report generation (Supplementary Table 10b). This highlights the value of the CT-RATE dataset and CT-CLIP in developing a generalist AI assistant for 3D medical imaging. Also, we demonstrate that engineering the system prompt of CT-CHAT (Supplementary Table 9) helps mitigate pathology hallucinations when the volume is not provided (Supplementary Figure 11). Our model, with new special tokens, can respond in various modes, including long and short answers, multiple-choice, and report generation (Extended Data Figure 9).
To further evaluate CT-CHAT’s clinical use, we benchmark its report generation performance against state-of-the-art 3D medical models RadFM [29] and CT2Rep [30]. Since radiology report generation closely mirrors the daily workflow of radiologists, it serves as a more clinically relevant benchmark than VQA. Instead of relying solely on conventional text similarity metrics, which are limited in assessing the clinical accuracy of generated content, we employ clinical efficacy metrics, using our in-house automated report labeler, which was also used to annotate CT-RATE (see creating the CT-RATE dataset section of Methods for more details). CT-CHAT consistently outperforms both RadFM and CT2Rep on the CT-RATE internal and RAD-ChestCT external validation sets across all clinical metrics (Figure 6e). Moreover, we demonstrate that integrating structured outputs from other advanced 3D algorithms can further improve report generation. By incorporating detailed nodule segmentation information, such as lobar location and size, extracted using a state-of-the-art 3D segmentation model, we substantially enhanced CT-CHAT’s report accuracy (Figure 6e). This experiment emphasizes that leveraging robust 3D models for different abnormalities could further boost performance, yet such high-quality models are still scarce. Continued development and open-source release of more domain-specific 3D segmentation algorithms are essential for future progress. Our expert evaluation adds a robust layer of validation, consistently confirming CT-CHAT’s superior performance (Figure 6d). This direct clinical assessment highlights the real-world value of our approach. Notably, CT-CHAT’s advantage over CT2Rep, despite both models being trained on the CT-RATE dataset, demonstrates the effectiveness of pretraining with CT-CLIP. These results reinforce the strength of our foundational model, CT-CLIP, in learning clinically meaningful representations and enabling high-quality performance in downstream tasks. That said, it remains far from ready for clinical deployment, as emphasized by our radiologists. As with all AI systems, challenges such as robustness, interpretability, regulatory compliance, and prospective validation must be addressed before integration into routine radiology workflows. To help broaden accessibility, we introduce a user-friendly graphical interface (via our GitHub repository) that allows non-experts to interact with the CT-CHAT model. This interface not only makes CT-CHAT a valuable tool for educational and clinical research settings but also invites a wider audience to engage with and contribute to the evolution of interactive AI-driven inference in clinical practice.
@@
Despite these advancements, our CT-RATE dataset and the CT-CLIP and CT-CHAT models have certain limitations. First, biases inherent to our single-institution dataset may influence our self-supervised training. Although we curate a large and diverse set of volume-report pairs, the data originates from a single geographic and clinical setting, which may limit the generalizability of our models to other populations. We conduct an ablation study to assess the impact of training set size on CT-CLIP’s performance, revealing that performance increases with larger training sets (Figure 4c). This suggests that incorporating data from multiple institutions and more diverse patient populations could further enhance model robustness. Second, CT-CLIP depends on the specific terminology used in reports and the choice of prompts during zero-shot detection. Variability in how diseases are described across radiology reports can affect detection accuracy. To assess the importance of prompt engineering, we evaluate zero-shot multi-abnormality detection using seven different prompts (Figure 4b). While we select the best-performing prompt from commonly used structures in reports, continued optimization may further improve performance. Third, the evaluation of CT-CHAT faces limitations due to the lack of pretrained 3D medical multimodal AI assistants for benchmarking. As a result, we compare CT-CHAT against 2D models, which may not fully capture the 3D spatial information inherent in CT volumes. To address this, we use digitally reconstructed radiographs (DRRs) generated from our dataset [46]. This approach is suggested by prior work showing that models trained with DRRs can yield comparable results to those trained with 3D chest CT volumes [46]. However, the absence of true 3D benchmarks highlights the need for further development and validation of multimodal AI assistants in 3D medical imaging. Due to the absence of 3D VQA baselines, we also benchmark CT-CHAT’s report generation function against state-of-the-art 3D radiology report generation models. Despite CT-CHAT being trained for VQA, CT-CHAT outperforms these models; however, the quality of its generated reports remains insufficient for clinical use, underscoring the need for further improvements in performance and clinical validation. Finally, our current models are limited to chest CT volumes. To broaden their clinical applicability, future work should focus on extending CT-RATE, CT-CLIP, and CT-CHAT to include other 3D imaging modalities such as MRI and PET scans, as well as other anatomical regions. Expanding to different modalities and body parts would enhance the use of our models across a wider range of medical imaging tasks and clinical scenarios.
In conclusion, the development of CT-CLIP and CT-CHAT marks a major step forward in 3D medical imaging. Enabled by CT-RATE, an open-source 3D dataset paired with radiology reports, these models set a benchmark for multimodal AI in this domain. Their open availability provides a strong foundation for future research and, ultimately, for the development of clinically deployable systems.
@@
Figure 1: Figure 1: Overview of our dataset and CT-CLIP. a. Data acquisition protocol for the CT-RATE dataset, comprising non-contrast 3D chest CT volumes and corresponding radiology text reports written by radiologists. b. Extraction of multi-abnormality labels from radiology reports by using a finetuned text encoder. c. Training of our CT-CLIP framework, which pairs 3D chest CT volumes with radiology reports from the CT-RATE dataset through a contrastive learning approach. d. Illustration of the contrastive learning procedure, employed in the correlation matrix during training to ensure precise match recognition and mismatch differentiation. $\text{V}_{\text{i}}$ shows volume embeddings and $\text{R}_{\text{i}}$ shows report embeddings. e. Inference for zero-shot multi-abnormality detection uses normalized probabilities derived from softmax-applied contrastive loss for each positive and negative prompt.
@@
Figure 2: Figure 2: Comprehensive analysis of the novel CT-RATE dataset. a. CT-RATE comprises data from 21,304 unique patients, with ages ranging from 18 to 102 years. b. For validation and finetuning purposes, multi-abnormality labels for 18 distinct abnormalities are extracted from the corresponding radiology reports for each CT volume. c. The training process of the CT-CLIP model uses the impression and findings sections of radiology reports, which vary in length, along with the chest CT volumes. d. The number of slices in the chest CT volumes within CT-RATE ranges from 100 to 600. e. Spacings in the x-axis, y-axis, and z-axis vary across volumes; all chest CT volumes are preprocessed to ensure consistent spacing before training each model. f. Despite originating from a single institution, CT-RATE includes volumes acquired using scanners from three different manufacturers, ensuring variability.
@@
Figure 4: Figure 4: A detailed evaluation of CT-CLIP a. Multi-abnormality classification performance is evaluated on an internal and two external validation sets using four key metrics. Our CT-CLIP outperforms the fully supervised baseline across all metrics. Performance improves with both finetuning techniques, demonstrating the robustness of the CT-CLIP’s encoder in feature extraction from CT volumes. For F1, a weighted score is computed per abnormality, then averaged across all findings to ensure balanced evaluation. The scores are color-highlighted to emphasize the performance increase compared to random benchmarks. b.Different prompts are used for zero-shot multi-abnormality detection with CT-CLIP. The best-performing prompt, highlighting the importance of prompt engineering, is selected for further experiments. Data are presented as mean ± SD of the bootstrapped values. c. To explore how the performance scales with dataset size, we train CT-CLIP using different portions of CT-RATE. (9.8% of CT-RATE matches the size of the only open-source CT dataset with multi-abnormality labels.) Contour lines denote $\pm$1 standard deviations calculated via bootstrapping, with the center line indicating the mean. d. t-SNE projections map the distribution of the number of abnormalities present in cases, using the latent values from individual reports and CT volumes. e. t-SNE projections illustrate the age distribution of patients. The clustering shows that CT volumes and reports from older patients tend to correspond with a higher number of abnormalities, indicating a correlation between age and disease prevalence.
@@
Figure 5: Figure 5: Using CT-CLIP for retrieval of 3D chest CT volumes a. CT-CLIP can be used for volume-to-volume retrieval using the vision encoder within CT-CLIP. b. CT-CLIP can be used for report-to-volume retrieval using both the vision and text transformers within CT-CLIP. c. Comprehensive evaluation of volume-to-volume retrieval performance on internal and external validation sets, with scores color-highlighted to emphasize improvements over random scores, shows similar retrieval accuracy even for the finetuned models. d. The evaluation of report-to-volume retrieval is limited to the internal validation set due to the absence of paired radiology reports in the external validation set. Notably, finetuning CT-CLIP for multi-abnormality classification diminishes its retrieval performance. This reduction in performance is primarily observed in report-to-volume retrieval, indicating that the VocabFine finetuning predominantly impacts on the text transformer component of CT-CLIP.
@@
Figure 6: Figure 6: Overview and detailed evaluation of CT-CHAT a. The architecture of CT-CHAT, a vision-language AI assistant for 3D chest CT volumes, trained by finetuning LoRA weights on a state-of-the-art large language model using CT tokens encoded by the 3D vision encoder of CT-CLIP. b. An illustration showing an example CT-CHAT output, highlighting its superior performance as a vision-language assistant specifically designed for 3D chest CT volumes compared to its 2D counterparts with DRRs provided. c. The question-answering performance of various vision-language assistants is evaluated across five key metrics, with radar charts providing a detailed breakdown of evaluations for each of the four different visual question-answering tasks. d. Report generation performance, evaluated by three board-certified radiologists and a violin plot of scores, confirms CT-CHAT’s clear superiority. e. Report generation performance is evaluated using clinical efficacy metrics, calculated by adapting CT-RATE’s automatic report labeler to assess clinical correctness across multiple findings on two test sets.